# SimpleMem: Efficient Lifelong Memory for LLM Agents

### v1

@@ Abstract
To support reliable long-term interaction in complex environments, LLM agents require memory systems that efficiently manage historical experiences. Existing approaches either retain full interaction histories via passive context extension, leading to substantial redundancy, or rely on iterative reasoning to filter noise, incurring high token costs. To address this challenge, we introduce SimpleMem, an efficient memory framework based on semantic lossless compression. We propose a three-stage pipeline designed to maximize information density and token utilization: (1) Semantic Structured Compression, which applies entropy-aware filtering to distill unstructured interactions into compact, multi-view indexed memory units; (2) Recursive Memory Consolidation, an asynchronous process that integrates related units into higher-level abstract representations to reduce redundancy; and (3) Adaptive Query-Aware Retrieval, which dynamically adjusts retrieval scope based on query complexity to construct precise context efficiently. Experiments on benchmark datasets show that our method consistently outperforms baseline approaches in accuracy, retrieval efficiency, and inference cost, achieving an average F1 improvement of 26.4% while reducing inference-time token consumption by up to 30×, demonstrating a superior balance between performance and efficiency. Code is available at https://github.com/aiming-lab/SimpleMem.
@@ 1 Introduction
Large Language Model (LLM) agents have recently demonstrated remarkable capabilities across a wide range of tasks (Xia et al., 2025; Team et al., 2025; Qiu et al., 2025). However, constrained by fixed context windows, existing agents exhibit significant limitations when engaging in long-context and multi-turn interaction scenarios (Liu et al., 2023; Wang et al., 2024a; Liu et al., 2025; Hu et al., 2025; Tu et al., 2025). To facilitate reliable long-term interaction, LLM agents require robust memory systems to efficiently manage and utilize historical experience (Dev & Taranjeet, 2024; Fang et al., 2025; Wang & Chen, 2025; Tang et al., 2025; Yang et al., 2025; Ouyang et al., 2025).
@@
Figure 1: Performance vs. Efficiency Trade-off. Comparison of Average F1 against Average Token Cost on the LoCoMo benchmark. SimpleMem occupies the ideal top-left position, achieving high accurac with minimal token consumption ($\sim$550 tokens).
To address these limitations, we introduce SimpleMem, an efficient memory framework inspired by the Complementary Learning Systems (CLS) theory (Kumaran et al., 2016) and designed around structured semantic compression. The core objective of SimpleMem is to improve information efficiency under fixed context and token budgets. To this end, we develop a three-stage pipeline that supports dynamic memory compression, organization, and adaptive retrieval: (1) Semantic Structured Compression: we apply an entropy-aware filtering mechanism that preserves information with high semantic utility while discarding redundant or low-value content. The retained information is reformulated into compact memory units and jointly indexed using dense semantic embeddings, sparse lexical features, and symbolic metadata, enabling multi-granular retrieval. (2) Recursive Memory Consolidation: Inspired by biological consolidation, we introduce an asynchronous process that incrementally reorganizes stored memory. Rather than accumulating episodic records verbatim, related memory units are recursively integrated into higher-level abstract representations, allowing repetitive or structurally similar experiences to be summarized while reducing semantic redundancy. (3) Adaptive Query-Aware Retrieval: we employ a query-aware retrieval strategy that dynamically adjusts retrieval scope based on estimated query complexity. Irrelevant candidates are pruned through lightweight symbolic and semantic constraints, enabling precise context construction tailored to task requirements. This adaptive mechanism achieves a favorable trade-off between reasoning performance and token efficiency.
@@ 2 The SimpleMem Architecture
In this section, we present SimpleMem, an efficient memory framework for LLM agents designed to improve information utilization under constrained context and token budgets through. As shown in Figure 2, the system operates through a three-stage pipeline. First, we describe Semantic Structured Compression process, which filters redundant interaction content and reformulates raw dialogue streams into compact memory units. Next, we describe Recursive Consolidation, an asynchronous process that incrementally integrates related memory units into higher-level abstract representations and maintaining a compact memory topology. Finally, we present Adaptive Query-Aware Retrieval, which dynamically adjusts retrieval scope based on estimated query complexity to construct precise and token-efficient contexts for downstream reasoning.
Figure 2: The SimpleMem Architecture. SimpleMem mitigates context inflation through three stages. (1) Semantic Structured Compression filters redundant interaction content and reformulates raw dialogue into compact, context-independent memory units. (2) Recursive Consolidation incrementally organizes related memory units into higher-level abstract representations, reducing redundancy in long-term memory. (3) Adaptive Query-Aware Retrieval dynamically adjusts retrieval scope based on query complexity, enabling efficient context construction under constrained token budgets.
@@ 2.1 Semantic Structured Compression
First, incoming dialogue is segmented into overlapping sliding windows $W_{t}$ of fixed length, where each window represents a short contiguous span of recent interaction. These windows serve as the basic units for evaluating whether new information should be stored. Then we employ a non-linear gating mechanism, $\Phi_{gate}$, to evaluate the information density of these dialogue windows to determine which windows is used fo indexing. For each window $W_{t}$, we compute an information score $H(W_{T})$ that jointly captures the introduction of new entities and semantic novelty relative to the immediate interaction history $H_{\text{prev}}$.
Formally, let $\mathcal{E}_{new}$ denote the set of named entities that appear in $W_{t}$ but not in $H_{\text{prev}}$. The information score is defined as:
$H(W_{t})=\alpha\cdot\frac{|\mathcal{E}_{new}|}{|W_{t}|}+(1-\alpha)\cdot(1-\cos(E(W_{t}),E(H_{prev})))$ (1)
where $E(\cdot)$ denotes a semantic embedding function and $\alpha$ controls the relative importance of entity-level novelty and semantic divergence.
Windows whose information score falls below threshold $\tau_{\text{redundant}}$ are treated as redundant and excluded from memory construction, meaning that the window is neither stored nor further processed, preventing low-utility interaction content from entering the memory buffer. For informative windows, the system proceeds to a segmentation step:
$\text{Action}(W_{t})=\begin{cases}\text{Segment}(W_{t}),&H(W_{t})\geq\tau_{\text{redundant}},\\ \varnothing,&\text{otherwise}.\end{cases}$ (2)
@@
For windows that pass the filter, we apply a segmentation function $\mathcal{F}_{\theta}$ to decompose each informative window into a set of context-independent memory units $m_{k}$}. This transformation resolves dependencies implicit in conversational flow by converting entangled dialogue into self-contained factual or event-level statements. Formally, $\mathcal{F}{\theta}$ is composed of a coreference resolution module ($\Phi_{\text{coref}}$) and a temporal anchoring module ($\Phi_{\text{time}}$):
$m_{k}=\mathcal{F}_{\theta}(W_{t})=\Phi_{time}\circ\Phi_{coref}\circ\Phi_{extract}(W_{t})$ (3)
@@
Here, $\Phi_{\text{extract}}$ identifies candidate factual statements, ($\Phi_{coref}$) replaces ambiguous pronouns with specific entity names (e.g., changing "He agreed" to "Bob agreed"), and $\Phi_{\text{time}}$ converts relative temporal expressions (e.g., transforming "next Friday" to "2025-10-24") into absolute ISO-8601 timestamps. This normalization ensures that each memory unit remains interpretable and valid independent of its original conversational context.
2.2 Structured Indexing and Recursive Consolidation
Then, the system need organize the resulting memory units to support efficient long-term storage and scalable retrieval. This stage consists of two components: (i) structured multi-view indexing for immediate access, and (ii) recursive consolidation for reducing redundancy and maintaining a compact memory topology over time.
@@ 2.2 Structured Indexing and Recursive Consolidation
To support flexible and precise retrieval, each memory unit is indexed through three complementary representations. First, at sematic layer, we map the entry to a dense vector space $\mathbf{v}_{k}$ using embedding models, which captures abstract meaning and enables fuzzy matching (e.g., retrieving "latte" when querying "hot drink"). Second, the Lexical Layer generates a sparse representation focusing on exact keyword matches and proper nouns, ensuring that specific entities are not diluted in vector space. Third, the Symbolic Layer extracts structured metadata, such as timestamps and entity types, to enable deterministic filtering logic. Formally, these projections form the comprehensive memory bank $\mathbb{M}$:
@@
$\mathbb{M}(m_{k})=\begin{cases}\mathbf{v}_{k}=E_{\text{dense}}(S_{k})\in\mathbb{R}^{d}&\text{(Semantic Layer)}\\ \mathbf{h}_{k}=\text{Sparse}(S_{k})\in\mathbb{R}^{|V|}&\text{(Lexical Layer)}\\ \mathcal{R}_{k}=\{(\text{key},\text{val})\}&\text{(Symbolic Layer)}\end{cases}$ (4)
@@
It allows the system to flexibly query information based on conceptual similarity, exact keyword matches, or structured metadata constraints.
While multi-view indexing supports efficient access, naively accumulating memory units over long interaction horizons leads to redundancy and fragmentation. To address this issue, we then introduces an asynchronous background consolidation process that incrementally reorganizes the memory topology. The consolidation mechanism identifies related memory units based on both semantic similarity and temporal proximity. For two memory units $m_{i}$ and $m_{j}$, we define an affinity score $\omega_{ij}$ as:
$\omega_{ij}=\beta\cdot\cos(\mathbf{v}_{i},\mathbf{v}_{j})+(1-\beta)\cdot e^{-\lambda|t_{i}-t_{j}|},$ (5)
where the first term captures semantic relatedness and the second term biases the model toward grouping events with strong temporal proximity.
When a group of memory units forms a dense cluster $\mathcal{C}$, determined by pairwise affinities exceeding a threshold $\tau_{\text{cluster}}$, the system performs a consolidation step:
$M_{\text{abs}}=\mathcal{G}_{\text{syn}}(\{m_{i}\mid m_{i}\in\mathcal{C}\}).$ (6)
This operation synthesizes repetitive or closely related memory units into a higher-level abstract representation $M_{\text{abs}}$, which captures their shared semantic structure. For example, instead of maintaining numerous individual records such as ‘‘the user ordered a latte at 8:00 AM,’’ the system consolidates them into a single abstract pattern, e.g., ‘‘the user regularly drinks coffee in the morning.’’ The original fine-grained entries are archived, reducing the active memory size while preserving the ability to recover detailed information if needed. As a result, the active memory index remains compact, and retrieval complexity scales gracefully with long-term interaction history.
2.3 Adaptive Query-Aware Retrieval
@@
After memory entries are organized, another challenge to retrieve relevant information efficiently under constrained context budgets. Standard retrieval approaches typically fetch a fixed number of context entries, which often results in either insufficient information or token wastage. To address this, we introduces an adaptive query-aware retrieval mechanism that dynamically adjusts retrieval scope based on estimated query complexity, thereby improving retrieval efficiency without sacrificing reasoning accuracy.
First, we propose a hybrid scoring function for information retrieval, $\mathcal{S}(q,m_{k})$, which aggregates signals from the tri-layer index established in the second stage. For a given query $q$, the relevance score is computed as:
$\mathcal{S}(q,m_{k})=$ $\lambda_{1}\cos(\mathbf{e}_{q},\mathbf{v}_{k})+\lambda_{2}\text{BM25}(q_{\text{lex}},S_{k})$ (7)
$+\gamma\,\mathbb{I}(\mathcal{R}_{k}\models\mathcal{C}_{\text{meta}}),$
where the first term measures semantic similarity in the dense embedding space, the second term captures exact lexical relevance, and the indicator function $\mathbb{I}(\cdot)$ enforces hard symbolic constraints such as entity-based filters.
Then, based on the hybrid scoring, we can rank the candidate memories by relevance. However, retrieving a fixed number of top-ranked entries remains inefficient when query demands vary. To address this, we estimate the query complexity $C_{q}\in[0,1]$, which reflects whether a query can be resolved via direct fact lookup or requires multi-step reasoning over multiple memory entries. A lightweight classifier predicts $C_{q}$ based on query features such as length, syntactic structure, and abstraction level.
$k_{dyn}=\lfloor k_{base}\cdot(1+\delta\cdot C_{q})\rfloor$ (8)
Based on this dynamic depth, the system modulates the retrieval scope. For low-complexity queries ($C_{q}\to 0$), the system retrieves only the top-$k_{min}$ high-level abstract memory entries or metadata summaries, minimizing token usage. Conversely, for high-complexity queries ($C_{q}\to 1$), it expands the scope to top-$k_{max}$, including a larger set of relevant entries, along with associated fine-grained details. The final context $\mathcal{C}_{final}$ is synthesized by concatenating these pruned results, ensuring high accuracy with minimal computational waste:
$\mathcal{C}_{final}=\bigoplus_{m\in\text{Top-}k_{dyn}(\mathcal{S})}[t_{m}:\text{Content}(m)]$ (9)
@@ 3 Experiments
In this section, we evaluate SimpleMem on the benchmark to answer the following research questions: (1) Does SimpleMem outperform other memory systems in complex long-term reasoning and temporal grounding tasks? (2) Can SimpleMem achieve a superior trade-off between retrieval accuracy and token consumption? (3) How effective are the proposed components? (4) What factors account for the observed performance and efficiency gains?

### v2

@@ Abstract
To support long-term interaction in complex environments, LLM agents require memory systems that manage historical experiences. Existing approaches either retain full interaction histories via passive context extension, leading to substantial redundancy, or rely on iterative reasoning to filter noise, incurring high token costs. To address this challenge, we introduce SimpleMem, an efficient memory framework based on semantic lossless compression. We propose a three-stage pipeline designed to maximize information density and token utilization: (1) Semantic Structured Compression, which distills unstructured interactions into compact, multi-view indexed memory units; (2) Online Semantic Synthesis, an intra-session process that instantly integrates related context into unified abstract representations to eliminate redundancy; and (3) Intent-Aware Retrieval Planning, which infers search intent to dynamically determine retrieval scope and construct precise context efficiently. Experiments on benchmark datasets show that our method consistently outperforms baseline approaches in accuracy, retrieval efficiency, and inference cost, achieving an average F1 improvement of 26.4% while reducing inference-time token consumption by up to 30×, demonstrating a superior balance between performance and efficiency. Code is available at https://github.com/aiming-lab/SimpleMem.
@@ 1 Introduction
Large Language Model (LLM) agents have recently demonstrated remarkable capabilities across a wide range of tasks (Xia et al., 2025; Team et al., 2025; Qiu et al., 2025). However, constrained by fixed context windows, existing agents exhibit significant limitations when engaging in long-context and multi-turn interaction scenarios (Liu et al., 2023; Wang et al., 2024a; Liu et al., 2025; Hu et al., 2025; Tu et al., 2025). To facilitate reliable long-term interaction, LLM agents require robust memory systems to efficiently manage and utilize historical experience (Dev and Taranjeet, 2024; Fang et al., 2025; Wang and Chen, 2025; Tang et al., 2025; Yang et al., 2025; Ouyang et al., 2025).
@@
Figure 1: Performance vs. Efficiency Trade-off. Comparison of F1 against Token Cost on the LoCoMo benchmark. SimpleMem achieves high accuracy with minimal token consumption.
To address these limitations, we introduce SimpleMem, an efficient memory framework inspired by the Complementary Learning Systems (CLS) theory (Kumaran et al., 2016) and built around structured semantic compression. The objective of SimpleMem is to improve information efficiency under fixed context and token budgets. We develop a three-stage pipeline that supports dynamic memory compression, organization, and adaptive retrieval: (1) Semantic Structured Compression: we apply a semantic density gating mechanism via LLM-based qualitative assessment. The system uses the foundation model as a semantic judge to estimate information gain relative to history, preserving only content with high downstream utility. Retained information is reformulated into compact memory units and indexed jointly using dense semantic embeddings, sparse lexical features, and symbolic metadata. (2) Online Semantic Synthesis: inspired by biological consolidation and optimized for real-time interaction, we introduce an intra-session process that reorganizes memory on-the-fly. Related memory units are synthesized into higher-level abstract representations during the write phase, allowing repetitive or structurally similar experiences to be denoised and compressed immediately. (3) Intent-Aware Retrieval Planning: we employ a planning-based retrieval strategy that infers latent search intent to determine retrieval scope dynamically. The system constructs a precise context by querying multiple indexes (symbolic, semantic, lexical) and unifying results through ID-based deduplication, balancing structural constraints and semantic relevance without complex linear weighting.
@@ 2 The SimpleMem Architecture
In this section, we present SimpleMem, which operates through a three-stage pipeline (see Figure 2 for the detailed architecture). Specifically, we first describe the Semantic Structured Compression, which utilizes implicit semantic gating to filter redundant interaction content and reformulate raw dialogue streams into compact memory units. Next, we describe Online Semantic Synthesis, an on-the-fly mechanism that instantly synthesizes related memory units into higher-level abstract representations, ensuring a compact and noise-free memory topology. Finally, we present Intent-Aware Retrieval Planning, which infers latent search intent to dynamically adjust retrieval scope, constructing precise and token-efficient contexts for downstream reasoning.
Figure 2: The SimpleMem Architecture. SimpleMem follows a three-stage pipeline: (1) Semantic Structured Compression filters low-utility dialogue and converts informative windows into compact, context-independent memory units. (2) Online Semantic Synthesis consolidates related fragments during writing, maintaining a compact and coherent memory topology. (3) Intent-Aware Retrieval Planning infers search intent to adapt retrieval scope and query forms, enabling parallel multi-view retrieval and token-efficient context construction.
@@ 2.1 Semantic Structured Compression
Specifically, first, incoming dialogue is segmented into overlapping sliding windows $W$ of fixed length, where each window represents a short contiguous span of recent interaction. These windows serve as the basic units for processing. Unlike traditional approaches that rely on rigid heuristic filters, we employ a Semantic Density Gating mechanism. We model the information assessment as an implicit reasoning task performed by the foundation model itself. The system leverages the attention mechanism of the LLM $f$ to estimate the information gain of the current window relative to the immediate history $H$.
Formally, we define the gating function $\Phi_{\text{gate}}$ as an indicator function that determines whether the semantic utility of $W$ exceeds the model’s implicit retention threshold $\tau$:
$\Phi_{\text{gate}}(W)=\mathbb{I}\left(f(W\mid H)>\tau\right).$ (1)
Windows where $\Phi_{\text{gate}}(W)=0$ are low-density (e.g., pure chitchat) and are discarded efficiently. This soft gating approach allows the system to capture subtle semantic nuances that rigid keyword or cosine-based filters often miss.
@@
For windows that pass the gate ($\Phi_{\text{gate}}(W)=1$), the system proceeds a normalization and decomposition step. We apply a transformation $\mathcal{F}_{\text{norm}}$ to decompose the window into a set of context-independent memory units $\{m_{k}\}_{k=1}^{K}$. This transformation resolves dependencies implicit in conversational flow by converting entangled dialogue into self-contained factual statements. Here, $\mathcal{F}_{\text{norm}}$ encapsulates three operations: fact extraction ($g_{\text{ext}}$), coreference resolution ($g_{\text{coref}}$), and temporal anchoring ($g_{\text{time}}$):
$m_{k}\in\mathcal{F}_{\text{norm}}(W)=g_{\text{time}}\circ g_{\text{coref}}\circ g_{ext}(W;f).$ (2)
@@
Here, $g_{\text{coref}}$ replaces ambiguous pronouns with specific entity names (e.g., changing "He agreed" to "Bob agreed"), and $g_{\text{time}}$ converts relative temporal expressions (e.g., transforming "next Friday" to "2025-10-24") into absolute ISO-8601 timestamps. This normalization ensures that each memory unit remains interpretable and valid independent of its original conversational context. By aggregating all resulting units $m_{k}$ across sliding windows, we obtain the complete memory set $\mathcal{M}$.
Following compression, the system organizes the memory units to support storage and retrieval. This stage consists of two synergistic processes: (i) structured multi-view indexing for precise access, and (ii) online semantic synthesis for minimizing redundancy at the point of creation.
@@
To support flexible and high-fidelity retrieval, each memory unit is indexed through three complementary representations. First, at the Semantic Layer, we map the entry to a dense vector space $s_{k}$ using embedding models, capturing abstract meaning to enable fuzzy matching (e.g., retrieving "latte" when querying "hot drink"). Second, the Lexical Layer utilizes an inverted index-based sparse representation. This acts as a high-dimensional sparse vector $l_{k}$ focusing on exact keyword matches and rare proper nouns, ensuring that specific entities are not diluted in dense vector space. Third, the Symbolic Layer extracts structured metadata, such as timestamps and entity types, to enable deterministic filtering logic. Formally, given a memory unit these projections form the comprehensive indexing $\mathcal{I}$:
@@
$\mathcal{I}(m_{t,k})=\begin{cases}s_{k}=E_{\text{dense}}(m_{k})&\text{(Semantic Layer)}\\ l_{k}=E_{\text{sparse}}(m_{k})&\text{(Lexical Layer)}\\ r_{k}=E_{\text{sym}}(m_{k})&\text{(Symbolic Layer)}\end{cases}$ (3)
@@
This architecture allows the system to flexibly query information based on conceptual similarity, exact keyword matches, or structured metadata constraints.
2.2 Online Semantic Synthesis
While this multi-view indexing strategy facilitates access, naively accumulating raw extractions leads to fragmentation, causing the memory structure to grow in a purely additive and unregulated manner that fails to adapt in real time to the evolving semantic context of an ongoing interaction. To address this, we introduce Online Semantic Synthesis, an intra-session consolidation mechanism. Unlike traditional systems that rely on asynchronous background maintenance, SimpleMem performs synthesis on-the-fly during the write phase. The model analyzes the stream of extracted facts within the current session scope and synthesizes related fragments into unified, high-density entries before they are committed to the database.
Formally, we define this synthesis as a transformation function $\mathcal{F}_{\text{syn}}$ that maps a set of new observations $O_{\text{session}}$ to a consolidated memory entry $\mathcal{F}_{\text{syn}}(O_{\text{session}},\mathcal{C}_{\text{context}};f)$, where $\mathcal{C}_{\text{context}}$ represents the current conversational context. This operation denoises the input by merging scattered details into a coherent whole. For instance, rather than storing three separate fragments like "User wants coffee", "User prefers oat milk", and "User likes it hot", the synthesis layer consolidates them into a single, comprehensive entry: "User prefers hot coffee with oat milk". This proactive synthesis ensures that the memory topology remains compact and free of redundant fragmentation, significantly reducing the burden on the retrieval system during future interactions.
2.3 Intent-Aware Retrieval Planning
@@
After memory entries are organized, the final challenge is to retrieve relevant information under constrained context budgets. Standard retrieval approaches typically fetch a fixed number of entries, which often results in recall failure for complex queries or token wastage for simple ones. To address this, we introduce Intent-Aware Retrieval Planning, a mechanism that dynamically determines the retrieval scope and depth by inferring the user’s latent search intent.
Unlike systems that rely on scalar complexity classifiers, SimpleMem leverages the reasoning capabilities of the LLM to generate a comprehensive retrieval plan. Given a query $q$ and history $H$, the planning module $\mathcal{P}$ acts as a reasoner to decompose the information needs and estimate the necessary search depth $d$:
$\{q_{\text{sem}},q_{\text{lex}},q_{\text{sym}},d\}\sim\mathcal{P}(q,H)$ (4)
where $q_{\text{sem}}$, $q_{\text{lex}}$, and $q_{\text{sym}}$ are optimized queries for semantic, lexical, and symbolic retrieval respectively. The parameter $d$ represents the adaptive retrieval depth, which reflects the estimated complexity of the query. Based on $d$, the system utilizes a candidate limit $n$ (where $n\propto d$) to balance recall coverage against context window constraints.
Guided by this plan, the system executes a parallel multi-view retrieval. We simultaneously query all three index layers defined in Section 2.1, imposing the quantity limit $n$ on each path:
$\mathcal{R}_{\text{sem}}$ $=\operatorname{Top-}n(\cos(E(q_{\text{sem}}),E(m_{i}))\mid m_{i}\in\mathcal{M})$ (5)
$\mathcal{R}_{\text{lex}}$ $=\operatorname{Top-}n(\operatorname{BM25}(q_{\text{lex}},m_{i})\mid m_{i}\in\mathcal{M})$
$\mathcal{R}_{\text{sym}}$ $=\operatorname{Top-}n(\{m_{i}\in\mathcal{M}\mid\text{Meta}(m_{i})\models q_{\text{sym}}\})$
Here, each view captures distinct relevance signals: $\mathcal{R}_{\text{sem}}$ retrieves based on dense embedding similarity; $\mathcal{R}_{\text{lex}}$ matches exact keywords or proper nouns; and $\mathcal{R}_{\text{sym}}$ filters entries based on structured metadata constraints.
Finally, we construct the context $\mathcal{C}_{q}$ by merging the results from these three views using a set union operation. This step naturally deduplicates overlapping entries, ensuring a comprehensive yet compact context:
$\mathcal{C}_{q}=\mathcal{R}_{\text{sem}}\cup\mathcal{R}_{\text{lex}}\cup\mathcal{R}_{\text{sym}}$ (6)
This hybrid approach ensures that strong signals from any view are preserved, allowing the system to adaptively scale its retrieval volume $n$ based on the inferred depth $d$.
@@ 3 Experiments
In this section, we evaluate SimpleMem on the benchmark to answer the following research questions: (1) Does SimpleMem outperform other memory systems in complex long-term reasoning understanding tasks? (2) Can SimpleMem achieve a superior trade-off between retrieval accuracy and token consumption? (3) How effective are the proposed components? (4) What factors account for the observed performance and efficiency gains?
@@
Table 2: Performance comparison on the LongMemEval benchmark. The evaluation uses gpt-4.1-mini as the judge. SimpleMem achieves the best overall performance while maintaining balanced capabilities across different sub-tasks.
Model Method Temporal Multi-Session Knowledge-Update Single-Session-User Single-Session-Assistant Single-Session-Preference Average
(n=133) (n=133) (n=78) (n=70) (n=56) (n=30)
GPT-4.1-mini Full-context 27.06% 30.08% 41.03% 47.14% 32.14% 60.00% 35.40%
Mem0 40.60% 50.37% 69.23% 87.14% 48.21% 63.33% 56.40%
LightMem 85.71% 47.37% 92.30% 88.57% 21.43% 76.67% 69.20%
SimpleMem 70.68% 60.92% 79.48% 85.71% 75.00% 76.67% 72.40%
GPT-4.1 Full-context 51.88% 39.10% 70.51% 65.71% 96.43% 16.67% 56.20%
Mem0 43.61% 54.89% 75.64% 54.29% 39.29% 83.33% 55.00%
LightMem 84.96% 57.89% 89.74% 87.14% 71.43% 70.00% 76.40%
SimpleMem 66.91% 81.20% 80.76% 98.57% 64.29% 76.78% 77.60%